How much energy (in joules) will be given off from the following reaction, if 3.0g of sodium reacts with 1.2g of nitrogen in the following reaction? 2Na+ N2 ——> 2NaN3 delta H= -42.7kJ
1
Expert's answer
2020-06-03T05:22:27-0400

In the reaction

2Na + 3N2 "\rightarrow" 2NaN3,

2 moles of sodium react with 3 moles of nitrogen:

"n(Na)\/2 = n(N_2)\/3" .

The number of the moles contained in 3 g of sodium is its mass divided by its molar mass "M" (22.99 g/mol):

"n(Na) = \frac{m(Na)}{M(Na)} = \frac{3.0}{22.99} = 0.1305" mol.

The number of the moles of nitrogen can also be calculated from its mass ("M_{N_2} =" 28.01 g/mol):

"n(N_2) = \frac{m(N_2)}{M(N_2)} = \frac{1.2}{28.01} = 0.04284" mol.

Let's compare the equivalent number of the moles of sodium and of nitrogen :

"3·n(Na)= 0.3915; 2·n(N_2) = 0.08568 \Rightarrow 3n(Na)>2n(N_2)" .

Therefore, there is an excess of sodium. It means that in order to calculate the energy released in the system, we must use the number of the moles of nitrogen. As 3 moles of nitrogen are used in the reaction, 3 moles of nitrogen are equivalent to the release of -42.7 kJ. So, the energy released is the product of the molar enthalpy and the number of the moles of nitrogen, divided by the conversion factor for nitrogen ( 3 mol N2) :

"Q = \u2206H\u00b7n(N_2)\/3 = 42.7\u00b70.04284\/3 = -0.610" kJ.

Answer: -0.610 kJ.